OpenStudy (anonymous):

r^2-d^2 factor completely.

jhonyy9 (jhonyy9):

@melissa1094 so do you know formula of difference of two squares ?

OpenStudy (anonymous):

no i dont /:

jhonyy9 (jhonyy9):

so can you tell me what is equal (a-b)(a+b) = ?

OpenStudy (anonymous):

would it be (a+b)^2

jhonyy9 (jhonyy9):

why ? (a+b)^2 =(a+b)(a+b) right ? but there i have wrote (a+b98a-b) yes ? so than ?

jhonyy9 (jhonyy9):

sorry there i have wrote (a+b)(a-b)

OpenStudy (anonymous):

so it wouldnt be (a+b)^2? all i know is that it has something to do with special product rules.

jhonyy9 (jhonyy9):

ok so you know that r*r-d*d=(r-d)(r+d)=r^2 +rd -rd -d^2 =r^2 -d^2

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

(r+d)(r-d)